According to Dietary Goals for the US (1977), high sodium intake may be related to ulcers, stomach cancer, and migraine headaches. The human requirement for salt is less than 220 milligrams per day, which is surpassed in most single servings of ready-to-eat noodles. If a random sample of 9 similar servings of noodles has a mean sodium content of 214 milligrams and a standard deviation of 7.5. What is the standard error of the 95% confidence interval of the mean sodium content of the noodles being investigated?



1
Expert's answer
2022-03-01T10:40:35-0500

The standard error is computed by dividing the sample standard deviation (s) by the square root of the sample size (n). That is:


Standard Error = "\frac{s}{\sqrt{n}}"


Given that s = 7.5 and n=9


The standard error = "\frac{7.5}{\sqrt{9}}" = 2.5


Answer: The standard error of the 95% confidence interval of the mean sodium content of the noodles being investigated is 2.5 milligrams